# Manuscript submitted to Aequitas 2024: Workshop on Fairness and Bias in AI | co-located with ECAI 2024, Santiago de Compostela, Spain

### v1

@@ Abstract
Motivated by the recital (67) of the current corrigendum of the AI Act in the European Union, we propose and present measures and mitigation strategies for discrimination in tabular datasets. We specifically focus on datasets that contain multiple protected attributes, such as nationality, age, and sex. This makes measuring and mitigating bias more challenging, as many existing methods are designed for a single protected attribute. This paper comes with a twofold contribution: Firstly, new discrimination measures are introduced. These measures are categorized in our framework along with existing ones, guiding researchers and practitioners in choosing the right measure to assess the fairness of the underlying dataset. Secondly, a novel application of an existing bias mitigation method, FairDo, is presented. We show that this strategy can mitigate any type of discrimination, including intersectional discrimination, by transforming the dataset. By conducting experiments on real-world datasets (Adult, Bank, Compas), we demonstrate that de-biasing datasets with multiple protected attributes is achievable. Further, the transformed fair datasets do not compromise any of the tested machine learning models’ performances significantly when trained on these datasets compared to the original datasets. Discrimination was reduced by up to 83% in our experimentation. For most experiments, the disparity between protected groups was reduced by at least 7% and 27% on average. Generally, the findings show that the mitigation strategy used is effective, and this study contributes to the ongoing discussion on the implementation of the European Union’s AI Act.
@@ 1 Introduction
Discrimination in artificial intelligence (AI) applications is a growing concern since the adoption of the AI Act by the European Parliament on March 13, 2024 [1]. It still remains a significant challenge across numerous domains [2, 3, 4, 5]. To prevent biased outcomes, pre-processing methods are often used to mitigate biases in datasets before training machine learning models [6, 7, 8, 9]. These methods are even mentioned in Recital (67) at the current corrigendum of the AI Act [1]:
@@
Since datasets often consist of multiple protected attributes, pre-processing methods should be able to handle these cases. However, only a few works have addressed it  [7, 10, 11, 12] and de-biasing datasets with multiple protected attributes is still an ongoing research topic.
Our paper mainly focuses on how to measure and mitigate discrimination in datasets where multiple protected attributes are present. In our first contribution, we provide a comprehensive categorization of discrimination measuring methods. Besides introducing new measures for some of these cases, we also categorize existing measures from the literature. Some of the listed measures specifically address intersectional discrimination and non-binary groups. The second contribution deals with bias mitigation. For this, we use an existing pre-processing framework, FairDo [9], that is fairness-agnostic. The fairness-agnostic property makes it possible to define any discrimination measure that should be minimized. By implementing the introduced measures, we can therefore mitigate biases for multiple protected attributes. Another advantage of FairDo is that it preserves data integrity and does not modify the features of individuals during the optimization process, unlike other methods [13, 3, 7].
We evaluated our methodology on popular tabular datasets with fairness concerns, such as Adult [14], Bank [15], and Compas [16]. These datasets contain multiple protected attributes and we mitigated the inherent biases. We used different discrimination measures to evaluate the effectiveness of the bias mitigation process. Because a successful mitigation process does not guarantee that the outcomes of machine learning models are fair, we trained machine learning models on the transformed datasets and evaluated their predictions regarding fairness and performance.
The results of the bias mitigation process as well as the performance of the machine learning models are promising. They indicate that achieving fairness in datasets with multiple protected attributes is perhaps possible, and FairDo is a proper framework for this task. Overall, our work contributes technical solutions for stakeholders to enhance the fairness of datasets and machine learning models, aiming for compliance with the AI Act [1].
Figure 1: Stick figures can be differentiated by their color and shape. In intersectional discrimination, attributes are intersected, which leads to new subgroups. In non-intersectional, each attribute is treated separately, i.e., colors and shapes are not intersecting in this case.
@@ 2 Preliminaries
To handle multiple protected attributes, we define $Z=\{Z_{1},\ldots,Z_{p}\}$ as a set of protected attributes. It can represent the set of sociodemographic features such as age, gender, and ethnicity. These factors may make individuals vulnerable to discrimination. Each protected attribute $Z_{k}\in Z$ is formally a discrete random variable that can take on values from a set $g_{k}$. In this context, we refer $g_{k}$ to groups that describe distinct social categories of a protected attribute. For example, let $Z_{k}$ represent gender; then $g_{k}$ is a set containing the genders male, female, and non-binary. To avoid limitations to a particular group fairness notion, we introduce a generalized notation based on the works of Žliobaitė [2], Duong and Conrad [9] in the following.
@@
With the definition of treatment, we can define fairness criteria that demand equal treatments for different groups. Let $P(E_{1}\mid E_{2},Z_{k}=i)$ and $P(E_{1}\mid E_{2},Z_{k}=j)$ be two treatments, then we call the following equation:
@@
To define statistical parity for the attribute $Z_{k}$ using our notation, we set $E_{1}\coloneqq(Y=1)$ and $E_{2}\coloneqq\text{True}$. By setting $E_{2}$ to a tautology, we compare the probabilities of the event $Y=1$ across different groups without conditioning on any additional event:
$P(Y=1\mid\text{True},Z_{k}=i)$ $=P(Y=1\mid\text{True},Z_{k}=j)$
@@
We use $\psi\colon\mathbb{D}\rightarrow\mathbb{R}$ to denote some discrimination measure that quantifies the discrimination inherent in any dataset $\mathcal{D}\in\mathbb{D}$. A dataset $\mathcal{D}$ consists of features, protected attributes, and labels for each individual. The explicit form of $\psi$ depends on cases introduced in Section 3.
@@ 3 Measuring Discrimination for Multiple Attributes
We found that numerous scenarios arise when dealing with multiple protected attributes. We categorize these scenarios based on the number of groups, denoted as $|g|$, and number of protected attributes, denoted as $|Z|$. By going through all cases, we present possible approaches from the literature as well as our own suggestions to measure discrimination.
3.1 Single Protected Attribute $(|Z|=1)$
In case of having only one protected attribute, i.e., $|Z|=|\{Z_{1}\}|=1$, we distinguish between cases by the number of available groups $|g|$ in the dataset. We categorize the cases by $|g|=0,1,2$, and $|g|>2$.
@@
When having no groups, the measurement of discrimination is impossible if no assumptions are being made. Discrimination can be assessed through proxy variables [19]; however, this approach can be imprecise and may introduce new biases. This case is equivalent to having no protected attribute, i.e., $|Z|=0$.
@@
    Difference to optimal treatment: Another way is to return the absolute difference of the group’s outcome to the optimal treatment. For example, group $i$ has an 80% chance of getting the favorable treatment. Ideally, having a 100% chance of receiving the favorable treatment would represent the optimal scenario. Therefore, the discrimination score is 20% in this case. It is given by:
@@
Without using any prior information, we can calculate the discrimination score by taking the absolute difference between the treatments of the two groups, as advised by Žliobaitė [2]. The discrimination measure $\psi$ is then simply given by the disparity as mentioned in Definition (2.3).
@@
While the case for binary attributes is straightforward, it becomes non-trivial for non-binary attributes that arise naturally in real-world data. We can fall back to $|g|=2$ by calculating the absolute difference between every distinct group $i,j\in g$. Because the discrimination between $i$ and $j$ is the same as between $j$ and $i$, only $\binom{|g|}{2}$ pairs need to be compared and we use an aggregation function $\operatorname{agg}^{(1)}$ to report the differences [2]. Lum et al. [21] refer to measures that aggregate or summarize discrimination scores as meta-metrics. The aggregate can be the sum or maximum function, depending on the use case. The result for a single protected attribute $Z_{k}$ with two or more groups can be computed as follows:
@@
where $\delta_{Z_{k}}$ is the disparity as defined in Definition (2.3) and $i<j$ ensures that each pair is considered only once (assuming label encoded groups). According to Žliobaitė [2] and her personal discussions with legal experts, she advocates using the maximum function, i.e.,
@@
A more general approach to measure discrimination is to calculate some form of correlation coefficient between the protected attribute and the outcome. The correlation coefficient can be calculated using Pearson’s correlation [22], Spearman or Kendall’s rank correlation [23, 24]. The discrimination measure can then be defined as the absolute value of the correlation coefficient:
@@
This approach can be applied to any number of groups. Fairlearn provides a pre-processing method that removes the correlation between the protected attribute and the outcome by transforming the data [7]. However, the given approach violates data integrity constraints as categorical attributes are transformed into continuous values. Further, zero correlation does not imply independency between two variables.
3.2 Multiple Protected Attributes $(|Z|>1)$
There are several ways to measure discrimination for multiple protected attributes ($|Z|>1$). Based on the works of Kearns et al. [20] and Kang et al. [12], we categorize them into two approaches: intersectional and non-intersectional (see Figure 1). Intersectional approaches consider the intersection of groups/identities. The overlapping of such identities forms subgroups [20]. Non-intersectional approaches treat each protected attribute independently.
@@ 3.2.1 Intersectional Discrimination
The central idea of intersectionality is that individuals experience overlapping forms of oppression or privilege based on the combination of multiple social categories they belong to. In the following, we will introduce necessary definitions to formulate intersectional discrimination, which is based on the work of Kearns et al. [20].
@@
Let $Z=\{Z_{1},\ldots,Z_{p}\}$ be a set of discrete random variables representing protected attributes that can take on values from corresponding groups $g_{1},\ldots,g_{p}$. A subgroup $i$ is defined as $i=(i_{1},\ldots,i_{p})\in g_{1}\times\ldots\times g_{p}$. In other words, a subgroup encompasses multiple groups from different protected attributes.
@@
Let $i$ be a subgroup as defined in Definition 3.1 and let $Z=\{Z_{1},\ldots,Z_{p}\}$ be a set of discrete random variables. Subgroup treatment is then defined as:
@@
Let $Z=\{Z_{1},\ldots,Z_{p}\}$ be a set of discrete random variables. Let $i,j\in g_{1}\times\ldots\times g_{p}$ be two subgroups with $i=(i_{1},\ldots,i_{p})$ and $j=(j_{1},\ldots,j_{p})$. The disparity between two subgroups is denoted as $\hat{\delta}_{Z}$ and is given by:
$\hat{\delta}_{Z}(i,j,E_{1},E_{2})=|P(E_{1}\mid E_{2},Z_{1}=i_{1},\ldots,Z_{p}=i_{p})-P(E_{1}\mid E_{2},Z_{1}=j_{1},\ldots,Z_{p}=j_{p})|.$
@@
Let denote the set of available subgroups as $G_{\text{avail}}\subseteq g_{1}\times\ldots\times g_{k}$. To finally capture the discrepancies across all available subgroup pairs, an aggregation function $\operatorname{agg}^{(1)}$ is applied to the subgroup disparities $\delta_{Z}$:
$\psi_{\text{intersect.}}(\mathcal{D})$ $=\underset{i,j\in G_{\text{avail}}}{\operatorname{agg}^{(1)}}\hat{\delta}_{Z}(i,j,E_{1},E_{2}).$ (8)
@@
$\psi_{\text{intersect.}}(\mathcal{D})$ $=\underset{i,j\in G_{\text{avail}}}{\max}\hat{\delta}_{Z_{k}}(i,j,E_{1},E_{2})$ (9)
$=\underset{i\in G_{\text{avail}}}{\max}\,P(E_{1}\mid E_{2},Z_{1}=i_{1},\ldots,Z_{p}=i_{p})-\underset{j\in G_{\text{avail}}}{\min}\,P(E_{1}\mid E_{2},Z_{1}=j_{1},\ldots,Z_{p}=j_{p}).$ (10)
Kang et al. [12] also dealt with intersectional discrimination in their work by introducing a multivariate random variable where each dimension represents a protected attribute. Their fairness objective is to minimize the mutual information between the outcome and the multivariate random variable. By minimizing the mutual information, the outcome is independent of the protected attributes, which is a desirable property for fairness [13, 25]. However, this approach relies on expensive techniques to approximate the mutual information. Using our notation, their formulation can be written as [12]:
@@
where MI denotes the mutual information and $Z$ is the multivariate random variable representing the protected attributes. We can trivially conclude that zero mutual information implies zero intersectional discrimination [12].

### v3

# AEQUITAS 2024: Workshop on Fairness and Bias in AI | co-located with ECAI 2024, Santiago de Compostela, Spain

@@ Abstract
Motivated by the recital (67) of the current corrigendum of the AI Act in the European Union, we propose and present measures and mitigation strategies for discrimination in tabular datasets. We specifically focus on datasets that contain multiple protected attributes, such as nationality, age, and sex. This makes measuring and mitigating bias more challenging, as many existing methods are designed for a single protected attribute. This paper comes with a twofold contribution: Firstly, new discrimination measures are introduced. These measures are categorized in our framework along with existing ones, guiding researchers and practitioners in choosing the right measure to assess the fairness of the underlying dataset. Secondly, a novel application of an existing bias mitigation method, FairDo, is presented. We show that this strategy can mitigate any type of discrimination, including intersectional discrimination, by transforming the dataset. By conducting experiments on real-world datasets (Adult, Bank, COMPAS), we demonstrate that de-biasing datasets with multiple protected attributes is possible. All transformed datasets show a reduction in discrimination, on average by 28%. Further, these datasets do not compromise any of the tested machine learning models’ performances significantly compared to the original datasets. Conclusively, this study demonstrates the effectiveness of the mitigation strategy used and contributes to the ongoing discussion on the implementation of the European Union’s AI Act.
@@ 1 Introduction
Discrimination in artificial intelligence (AI) applications is a growing concern since the adoption of the AI Act by the European Parliament on March 13, 2024 [1]. It still remains a significant challenge across numerous domains [2, 3, 4, 5]. To prevent biased outcomes, pre-processing methods are often used to mitigate biases in datasets before training machine learning models [6, 7, 8, 9]. The current corrigendum of the AI Act [1] emphasizes this in Recital (67):
@@
Since datasets often consist of multiple protected attributes, pre-processing methods should be able to handle these cases. However, only a few works have addressed this issue [7, 10, 11, 12, 13] and de-biasing such datasets is still an ongoing research topic. In addition, there is no straightforward approach to managing multiple protected attributes, as shown in Figure 1.
Our paper mainly focuses on how to measure and mitigate discrimination in datasets where multiple protected attributes are present. In our first contribution, we provide a comprehensive categorization of discrimination measuring methods. Besides introducing new measures for some of these cases, we also categorize existing measures from the literature. Some of the listed measures specifically address intersectional discrimination and non-binary groups. The second contribution deals with bias mitigation. For this, we use our published pre-processing framework, FairDo [9], that is fairness-agnostic. The fairness-agnostic property makes it possible to define any discrimination measure that should be minimized. By implementing the introduced measures, we can therefore mitigate biases for multiple protected attributes. Another advantage of FairDo is that it preserves data integrity and does not modify the features of individuals during the optimization process, unlike other methods [14, 3, 7].
We evaluated our methodology on popular tabular datasets with fairness concerns, such as Adult [15], Bank [16], and COMPAS [17]. We used different discrimination measures to evaluate the effectiveness of the bias mitigation process. Because a successful mitigation process does not guarantee that the outcomes of machine learning models are fair, we trained machine learning models on the transformed datasets and evaluated their predictions regarding fairness and performance. The code for the experiments can be found in the accompanying repository: https://github.com/mkduong-ai/fairdo/evaluation.
The results of the bias mitigation process as well as the performance of the machine learning models are promising. They indicate that achieving fairness in datasets with multiple protected attributes is possible, and FairDo is a proper framework for this task. Overall, our work contributes technical solutions for stakeholders to enhance the fairness of datasets and machine learning models, aiming for compliance with the AI Act [1].
Figure 1: Stick figures can be differentiated by their color and shape. In intersectional discrimination, attributes are intersected, which leads to new subgroups. In non-intersectional, each attribute is treated independently, i.e., colors and shapes are not intersecting in this case.
@@ 2 Preliminaries
To handle multiple protected attributes, we define $\mathcal{Z}=\{Z_{1},\ldots,Z_{p}\}$ as a set of protected attributes. It can represent the set of sociodemographic features such as age, gender, and ethnicity. These factors may make individuals vulnerable to discrimination. Each protected attribute $Z_{k}\in\mathcal{Z}$ is formally a discrete random variable that can take on values from the sample space $g_{k}$. In this context, we refer $g_{k}$ to groups that describe distinct social categories of a protected attribute. For example, let $Z_{k}$ represent gender; then $g_{k}$ is a set containing the genders male, female, and non-binary. To avoid limitations to a particular group fairness notion, we introduce a generalized notation based on the works of Žliobaitė [2], Duong and Conrad [9] in the following.
@@
With the definition of treatment, we can define fairness criteria that demand equal treatment for different groups. Let $P(E_{1}\mid E_{2},Z_{k}=i)$ and $P(E_{1}\mid E_{2},Z_{k}=j)$ be treatments, then we call the following equation:
@@
To define statistical parity for the attribute $Z_{k}$ using our notation, we set $E_{1}\coloneqq(Y=1)$ and $E_{2}\coloneqq\Omega$. By setting $E_{2}$ to the sample space $\Omega$, we compare the probabilities of the event $Y=1$ across different groups without conditioning on any additional event:
$P(Y=1\mid\Omega,Z_{k}=i)$ $=P(Y=1\mid\Omega,Z_{k}=j)$
@@
We use $\psi\colon\mathbb{D}\rightarrow\mathbb{R}$ to denote some discrimination measure that quantifies the discrimination inherent in any dataset $\mathcal{D}\in\mathbb{D}$. A dataset $\mathcal{D}$ consists of features, protected attributes, and labels for each individual. The explicit form of $\psi$ depends on the cases introduced in Section 3.
@@ 3 Measuring Discrimination for Multiple Attributes
We found that numerous scenarios arise when dealing with multiple protected attributes. We categorize these scenarios based on the number of groups, denoted as $|g|$, and the number of protected attributes, denoted as $|\mathcal{Z}|$. By going through all cases, we present possible approaches from the literature as well as our own suggestions to measure discrimination.
3.1 Single Protected Attribute $(|\mathcal{Z}|=1)$
In the case of having only one protected attribute, i.e., $|\mathcal{Z}|=|\{Z_{1}\}|=1$, we distinguish between cases by the number of available groups $|g|$ in the dataset. We categorize the cases by $|g|=0,1,2$, and $|g|>2$.
@@
When there are no groups, the measurement of discrimination is impossible if no assumptions are being made. Discrimination can be assessed through proxy variables [20]; however, this approach can be imprecise and may introduce new biases. This case is equivalent to having no protected attribute, i.e., $|\mathcal{Z}|=0$.
@@
    Difference to optimal treatment: Another way is to return the absolute difference of the group’s outcome to the optimal treatment. For example, group $i$ has an 80% chance of receiving the favorable treatment. Ideally, having a 100% chance would represent the optimal scenario. Therefore, the discrimination score is 20% in this case. It is given by:
@@
Without using any prior information, we can calculate the discrimination score by taking the absolute difference between the treatments of the two groups, as advised by Žliobaitė [2]. The discrimination measure $\psi$ is then simply given by the disparity as mentioned in Definition 2.3.
@@
While the case for binary attributes is straightforward, it becomes non-trivial for non-binary attributes that arise naturally in real-world data. We can fall back to $|g|=2$ by calculating the absolute difference between every distinct group $i,j\in g$. Because the discrimination between $i$ and $j$ is the same as between $j$ and $i$, only $\binom{|g|}{2}$ pairs need to be compared and we use an aggregation function $\operatorname{agg}^{(1)}$ to report the differences [2]. Lum et al. [22] refers to measures that aggregate or summarize discrimination scores as meta-metrics. The aggregate can be the sum or maximum function, depending on the use case. The result for a single protected attribute $Z_{k}$ with two or more groups can be computed as follows:
@@
where $\delta_{Z_{k}}$ is the disparity as defined in Definition 2.3 and $i<j$ ensures that each pair is considered only once (assuming label-encoded groups). According to Žliobaitė [2] and her personal discussions with legal experts, she advocates using the maximum function, i.e.,
@@
A more general approach to measuring discrimination is to calculate some form of correlation coefficient between the protected attribute and the outcome. The correlation coefficient can be calculated using Pearson’s correlation [23], Spearman or Kendall’s rank correlation [24, 25]. The discrimination measure can then be defined as the absolute value of the correlation coefficient:
@@
This approach can be applied to any number of groups. Fairlearn provides a pre-processing method that removes the correlation between the protected attribute and the outcome by transforming the data [7]. However, the given approach violates data integrity constraints as categorical attributes are transformed into continuous values. Moreover, zero correlation does not imply independence between two variables.
3.2 Multiple Protected Attributes $(|\mathcal{Z}|>1)$
There are several ways to measure discrimination for multiple protected attributes ($|\mathcal{Z}|>1$). Based on the works of Kearns et al. [21], Yang et al. [11] and Kang et al. [13], we categorize them into two approaches: intersectional and non-intersectional (see Figure 1). Intersectional approaches consider the intersection of identities. The overlapping of such identities forms subgroups [21]. Non-intersectional approaches treat each protected attribute independently [11].
@@ 3.2.1 Intersectional Discrimination
The central idea of intersectionality is that individuals experience overlapping forms of oppression or privilege based on the combination of multiple social categories they belong to. In the following, we will introduce definitions to formulate intersectional discrimination, which is based on the work of Kearns et al. [21].
@@
Let $\mathcal{Z}=\{Z_{1},\ldots,Z_{p}\}$ be a set of discrete random variables representing protected attributes that can take on values from corresponding groups $g_{1},\ldots,g_{p}$. A subgroup $i$ is defined as $i=(i_{1},\ldots,i_{p})\in g_{1}\times\ldots\times g_{p}$. In other words, a subgroup encompasses multiple groups from different protected attributes.
@@
Let $i$ be a subgroup as defined in Definition 3.1 and let $\mathcal{Z}=\{Z_{1},\ldots,Z_{p}\}$ be a set of discrete random variables. Subgroup treatment is then defined as:
@@
Let $\mathcal{Z}=\{Z_{1},\ldots,Z_{p}\}$ be a set of discrete random variables. Let $i,j\in g_{1}\times\ldots\times g_{p}$ be two subgroups with $i=(i_{1},\ldots,i_{p})$ and $j=(j_{1},\ldots,j_{p})$. The disparity between two subgroups is denoted as $\hat{\delta}_{\mathcal{Z}}$ and is given by:
$\hat{\delta}_{\mathcal{Z}}(i,j,E_{1},E_{2})=|P(E_{1}\mid E_{2},Z_{1}=i_{1},\ldots,Z_{p}=i_{p})-P(E_{1}\mid E_{2},Z_{1}=j_{1},\ldots,Z_{p}=j_{p})|.$
@@
Let us denote the set of available subgroups as $G_{\text{avail}}\subseteq g_{1}\times\ldots\times g_{k}$. To finally capture the discrepancies across all available subgroup pairs, an aggregation function $\operatorname{agg}^{(1)}$ is applied to the subgroup disparities $\hat{\delta}_{\mathcal{Z}}$:
$\psi_{\text{intersect}}(\mathcal{D})$ $=\underset{i,j\in G_{\text{avail}}}{\operatorname{agg}^{(1)}}\hat{\delta}_{\mathcal{Z}}(i,j,E_{1},E_{2}).$ (8)
@@
$\psi_{\text{intersect}}(\mathcal{D})$ $=\underset{i,j\in G_{\text{avail}}}{\max}\hat{\delta}_{Z_{k}}(i,j,E_{1},E_{2})$ (9)
$=\underset{i\in G_{\text{avail}}}{\max}\,P(E_{1}\mid E_{2},Z_{1}=i_{1},\ldots,Z_{p}=i_{p})-\underset{j\in G_{\text{avail}}}{\min}\,P(E_{1}\mid E_{2},Z_{1}=j_{1},\ldots,Z_{p}=j_{p}).$
Kang et al. [13] also dealt with intersectional discrimination in their work by introducing a multivariate random variable $Z$ where each dimension represents a protected attribute. Their fairness objective is to minimize the mutual information between the outcome and the multivariate random variable. By minimizing the mutual information, the outcome is independent of the protected attributes, which is a desirable property for fairness [14, 26]. In this context, zero mutual information implies the absence of intersectional discrimination [13]. However, this approach relies on expensive techniques to approximate the mutual information. Using our notation, their formulation can be written as [13]:
@@
where MI denotes the mutual information.